# Quantum Non-Local Nonstabilizerness

### v3

@@ Abstract
Quantum entanglement and quantum nonstabilizerness are fundamental resources that characterize distinct aspects of a quantum state: entanglement reflects non-local correlations, while nonstabilizerness quantifies the deviation from stabilizer states. A quantum state becomes a valuable resource for applications like universal quantum computation only when both quantities are present. Here, we propose that quantum non-local nonstabilizerness (NN) serves as an effective measure of this combined resource, incorporating both entanglement and nonstabilizerness. We demonstrate that NN can be precisely computed for two-qubit pure states, where it is directly related to the entanglement spectrum. We then extend the definition of NN to mixed states and explore its presence in many-body quantum systems, revealing that the two-point NN decays according to a power law in critical states. Furthermore, we explore measurement-induced NN and uncover an intriguing phenomenon termed “nonstabilizerness swapping”, analogous to entanglement swapping, wherein post-measurement NN decays more slowly than any pre-measurement correlations. Our results thus represent a pivotal step towards accurately quantifying the “quantumness” of a state and reveal the potential for manipulating this resource through measurements.
@@
Quantum entanglement, which captures the inability of a quantum state to be decomposed into product states of its subsystems, is a fundamental manifestation of quantum mechanics [1]. It serves as a cornerstone for numerous applications, including quantum computation, quantum simulation, and quantum communication [2, 3]. Entanglement is conventionally viewed as a non-local resource that cannot be generated through local operations alone, and is typically quantified by the von Neumann entropy for pure states or the entanglement negativity for mixed states [4, 5, 6, 7, 8]. Since an entangled state of $N$ qubits generally requires storing $O(2^{N})$ complex amplitudes for its classical representation, quantum algorithms leveraging such states can achieve exponential speedups over direct classical simulations, potentially even surpassing any known classical algorithm—a concept referred to as quantum supremacy [9, 10, 11].
@@
Nevertheless, entanglement alone does not fully capture the “quantumness” of a state. Certain highly entangled states, known as stabilizer states, can still be efficiently described and simulated [12, 13, 14]. A stabilizer state is uniquely specified by a set of stabilizer generators, which requires only $O(N^{2})$ memory space. Furthermore, when a stabilizer state evolves under unitary gates from the Clifford group or under Pauli measurements, it remains within the stabilizer formalism, allowing for efficient classical simulation of the entire process. Consequently, nonstabilizerness, also referred to as “magic”, is equally crucial for a quantum state to serve as a valuable resource in demonstrating quantum supremacy [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. It is worth emphasizing that a key difference between these two resources is that entanglement remains invariant under local unitary transformations, whereas nonstabilizerness is basis-dependent, implying that it can be altered through local operations.
@@
Thus, a natural question arises: Is there a suitable measure of the genuinely non-classical resources that integrates both entanglement and nonstabilizerness? Moreover, is this measure computationally tractable? Theoretically, such a measure would deepen our understanding of the connection between these two seemingly distinct properties [35, 36], while practically it would help characterize algorithmic complexity and guide the selection of initial states in quantum computations and simulations. With the existence of such a measure, it becomes pertinent to identify resourceful states [37]. Furthermore, a more intriguing question is whether this resource can be manipulated through measurement.
@@
In this Letter, we address these questions by investigating quantum non-local nonstabilizerness (NN), which quantifies the minimal amount of nonstabilizerness contained in a bipartite state after optimizing over all possible local unitary transformations on each subsystem [38, 39, 40]. We first demonstrate that NN is a promising candidate for characterizing quantum resources and is applicable to any faithful and additive nonstabilizerness measure. Although previous studies have focused on providing bounds and estimates of NN, we show that NN is analytically tractable for a two-qubit pure state and directly related to the entanglement spectrum. We then investigate which states exhibit nonzero NN. Our results indicate that NN is ubiquitous in typical two-qubit pure states, with Haar-random states having the highest probability of attaining maximal NN. Extending the definition to the mixed state, we further examine the behavior of two-point NN in critical many-body states, using both the transverse-field Ising model (TFIM) at the critical point and a monitored Haar-random unitary circuit (MHC) with critical measurement rate as representative examples [41, 42, 43]. We find that two-point NN decays following a power law, mirroring the behavior of mutual information. Finally, we explore whether NN can be manipulated through measurements, akin to the manipulation of entanglement. Intriguingly, measurement-induced NN in a critical MHC decays more slowly than any correlation present before measurement, revealing a “nonstabilizerness swapping” phenomenon reminiscent of entanglement swapping [44, 45].
@@
Non-local nonstabilizerness. We consider a bipartite pure state $\left|\Psi_{AB}\right\rangle$, where $A$ and $B$ represent two subsystems. As illustrated in Fig. 1(a), the NN is defined as [38, 39]:
@@
where $M$ is a measure of nonstabilizerness, and $U_{A}$ and $U_{B}$ are unitary gates acting on subsystems $A$ and $B$, respectively. We will omit the subscript of $\mathcal{M}_{AB}$ when the bipartition is clear from the context. In operational terms, a finite NN indicates that the state cannot be efficiently represented as a stabilizer state, regardless of the local unitaries applied to subsystems $A$ and $B$ in an effort to reduce the nonstabilizerness. By definition, $\mathcal{M}$ is symmetric under the exchange of subsystems $A$ and $B$ and is stable under local basis rotations, which can be regarded as free operations. Since $\mathcal{M}$ can be directly related to the flatness of the entanglement spectrum, a nonzero value of $\mathcal{M}$ directly implies that the state possesses both entanglement and nonstabilizerness [46]. For any faithful nonstabilizerness measure $M$, $\mathcal{M}$ would be further guaranteed to be non-negative. Moreover, if $M$ is additive, i.e., $M(\left|\psi\right\rangle\otimes\left|\phi\right\rangle)=M(\left|\psi\right\rangle)+M(\left|\phi\right\rangle)$, then the NN is sub-additive. Specifically, for two bipartite pure states $\left|\Psi_{A_{1}B_{1}}\right\rangle$ and $\left|\Phi_{A_{2}B_{2}}\right\rangle$, we have $\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle)\geq\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. The proof can be found in [47]. In summary, for any additive and faithful nonstabilizerness measure, the NN is stable, sub-additive, symmetric and vanishes for both product states and stabilizer states. Given these desirable properties and clear physical significance, $\mathcal{M}$ serves as an effective measure to characterize the novel resources of a quantum state that account for both nonstabilizerness and entanglement.
@@
Although various nonstabilizerness measures satisfy faithfulness and additivity, we select $M$ to be the second stabilizer Rényi entropy (SRE) in the following, as it is well defined for qubits and avoids any minimization procedure in its calculation [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. It is given by:
$M(\rho)=-\text{log}\left(\frac{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{4}}{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{2}}\right),$ (2)
@@
which for pure states reduces to $M(\left|\psi\right\rangle)=-\text{log}\sum_{P\in\mathbb{P}_{N}}(1/2^{N})|\left\langle\psi|P|\psi\right\rangle|^{4}$, where $N$ is the number of qubits and $\mathbb{P}_{N}$ represents the set of $N$-fold tensor products of Pauli matrices, denoted as $\sigma_{\mu}$ with $\mu=0,x,y,z$. It is also instructive to compare $\mathcal{M}$ with the mutual SRE introduced in previous work, which is defined as $I_{AB}(\left|\Psi_{AB}\right\rangle)=M(\left|\Psi_{AB}\right\rangle)-M(\rho_{A})-M(\rho_{B})$, where $\rho_{A}$ and $\rho_{B}$ are reduced states of subsystems $A$ and $B$, respectively [59, 61, 60]. While the mutual SRE captures certain aspects of the non-locality of a state’s nonstabilizerness, it lacks stability under local unitary transformations [65, 66], can take negative values in some cases, and generally does not align with the NN $\mathcal{M}$, as we will demonstrate later.
@@
Two-qubit NN. In general, computing $\mathcal{M}$ is highly complicated because it involves minimizing over the entire spectrum of local unitary transformations. However, for the two-qubit case, we demonstrate that an analytic expression for $\mathcal{M}$ can be derived. We first prove the following theorem.
@@
Theorem: For a two-qubit pure state of the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$, $\mathcal{M}(\left|\psi\right\rangle)=M(\left|\psi\right\rangle)=\text{log}(\frac{8}{7+\text{cos}(8\theta)})$.
@@
The proof is provided in [47], where the key idea is that the Pauli basis representation of this specific state is in a canonical form, with the reduced density matrices of both qubits exhibiting fully polarized Bloch vectors, and the full state having a diagonal correlation matrix [67, 68]. The following important corollary arises:
@@
Corollary: For any two-qubit pure state $\left|\Psi\right\rangle$ with the bipartite entanglement spectrum $\{\text{cos}^{2}(\theta),\text{sin}^{2}(\theta)\}$, $\mathcal{M}(\left|\Psi\right\rangle)=\text{log}(\frac{8}{7+\text{cos}(8\theta)})$.
@@
This holds because any two-qubit pure state can be transformed into the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$ by applying an appropriate local basis rotation, which can be determined via singular value decomposition. As a result, the NN is directly related to the entanglement spectrum and can be efficiently computed. It should be noted that NN has previously been understood to be approximated by quantities derived from the entanglement spectrum [38]. Our results show that this approximation is, in fact, exact for the two-qubit case. A direct implication is that any two-qubit pure state can have at most $\text{log}(4/3)$ amount of NN, which equals the nonstabilizerness of the H-type magic state [19]. Additionally, it is noteworthy that, when combined with the sub-additivity of $\mathcal{M}$, if a many-body state can be decomposed into product states of multiple two-qubit states of the above form, an upper bound for its NN can be easily determined.
@@
Subsequently, we investigate the typical amount of NN in a random two-qubit pure state. We generate random pure states via a Haar-random two-qubit unitary gate and examine the distribution of $\mathcal{M}$ across these states. The results, shown in Fig. 1(c), indicate that the state most commonly hosts the maximal $\mathcal{M}$. While the analytical expression for this probability distribution can be derived [47], a closed-form expression for the average value of $\mathcal{M}$, analogous to that for entanglement [69, 70], remains unclear. Nevertheless, numerical results suggest that the average $\mathcal{M}$ is approximately $0.192$.
@@
Mixed state generalization. For mixed states, the NN can be defined analogously to Eq. (1) as $\mathcal{M}_{AB}(\rho_{AB})=\min_{U_{A}\otimes U_{B}}M((U_{A}\otimes U_{B})\rho_{AB}(U_{A}\otimes U_{B})^{\dagger})$. Under this definition, the NN retains the properties of stability, symmetry and subadditivity for any additive measure of nonstabilizerness. By continuing to employ the second SRE, which is faithful when the mixed stabilizer states are regarded resource-free, the NN also vanishes for the mixed stabilizer states [48, 71]. However, a crucial difference is that the NN does not vanish for certain separable states. This difference arises because, even for a single-qubit mixed state, such as $\rho=\frac{1}{2}(\sigma_{0}+\frac{1}{2}\sigma_{x})$, no unitary transformation can bring it back to a mixed stabilizer state. We also explore other possible nonstabilizerness measures for the mixed state, while the conclusion remains the same [47].
Focusing once again on the two-qubit case, it follows directly from the proof of the previous theorem that if a two-qubit mixed state $\rho$ has a Pauli basis representation in canonical form, we would have $\mathcal{M}(\rho)=M(\rho)$. However, it is important to note that not all mixed states can be transformed into the canonical form using only local unitary transformations [67]. In such cases, numerical optimization is inevitable, as examples will be shown in the next section.
@@
Consider the Werner state as an analytically tractable example, which is defined as $\rho_{W}=x\left|\psi_{-}\right\rangle\left\langle\psi_{-}\right|+(1-x)\mathbb{I}/4$, with $|\psi_{-}\rangle\equiv(|01\rangle-|10\rangle)/\sqrt{2}$ [72]. This state has vanishing Bloch vectors for each qubit and a diagonal correlation matrix $T$ with diagonal entries given by $-x/4$. We calculate NN and logarithmic negativity $E_{N}$, a measure of separability, with the results shown in Fig. 1(d). Although the Werner state remains separable for $x<1/3$, the NN takes on a finite value for small $x$. As $x$ approaches $1$, the state becomes more entangled, yet the NN approaches zero as the state becomes a stabilizer state. The difference between $\mathcal{M}$ and total nonstabilizerness $M$ again becomes evident if we further apply a T-gate on the first qubit. While the NN remain unaffected, the total nonstabilizerness $M$ would change significantly.

### v4

@@ Abstract
Quantum entanglement and quantum nonstabilizerness are fundamental resources that characterize distinct aspects of a quantum state: entanglement reflects non-local correlations, while nonstabilizerness quantifies the deviation from stabilizer states. A quantum state becomes a valuable resource for applications like universal quantum computation only when both quantities are present. Here, we propose that quantum non-local nonstabilizerness (NN) serves as an effective measure of this combined resource, incorporating both entanglement and nonstabilizerness. We demonstrate that NN can be precisely computed for two-qubit pure states, establishing a direct relationship with the entanglement spectrum. We then extend the definition of NN to mixed states and explore its presence in many-body quantum systems. We reveal that the two-point NN decays according to a power law in critical states, indicating that it can serve as an effective quantum information probe for phase transitions. Furthermore, we explore measurement-induced NN and uncover an intriguing phenomenon termed “nonstabilizerness swapping”, wherein post-measurement NN decays more slowly than any pre-measurement correlations. The occurrence of this phenomenon provides a sufficient condition for both entanglement swapping and the presence of a nontrivial sign structure. Our results thus represent a pivotal step toward accurately quantifying the “quantumness” of a state and bring insights from multiple perspectives into the property of NN.
I Introduction
@@
Quantum entanglement, which captures the inability of a quantum state to be decomposed into product states of its subsystems, is a fundamental manifestation of quantum mechanics Horodecki et al. (2009). It serves as a cornerstone for numerous applications, including quantum computation, quantum simulation, and quantum communication Nielsen and Chuang (2010); Wilde (2017). Entanglement is conventionally viewed as a non-local resource that cannot be generated through local operations alone, and is typically quantified by the von Neumann entropy for pure states or the entanglement negativity for mixed states Bennett et al. (1996); Plenio (2005); Vidal and Werner (2002); Życzkowski et al. (1998); Horodecki and Horodecki (1996). Since an entangled state of $N$ qubits generally requires storing $O(2^{N})$ complex amplitudes for its classical representation, quantum algorithms leveraging such states can achieve exponential speedups over direct classical simulations, potentially even surpassing any known classical algorithm—a concept referred to as quantum supremacy or, more recently, quantum advantage Arute et al. (2019); Boixo et al. (2018); Harrow and Montanaro (2017).
@@
Nevertheless, entanglement alone does not fully capture the “quantumness” of a state. Certain highly entangled states, known as stabilizer states, can still be efficiently described and simulated Gottesman (1998, 1997); Aaronson and Gottesman (2004). A stabilizer state is uniquely specified by a set of stabilizer generators, which requires only $O(N^{2})$ memory space. Furthermore, when a stabilizer state evolves under unitary gates from the Clifford group or under Pauli measurements, it remains within the stabilizer formalism, allowing for efficient classical simulation of the entire process. Consequently, nonstabilizerness, also referred to as “magic”, is equally crucial for a quantum state to serve as a valuable resource in demonstrating quantum advantage Beverland et al. (2020); Bravyi and Gosset (2016); Bravyi et al. (2019, 2016); Bravyi and Kitaev (2005); Bu and Koh (2019); Campbell (2011); Gu et al. (2024a, 2025, b); Howard and Campbell (2017); Liu and Winter (2022); Seddon et al. (2021); True and Hamma (2022); Yoganathan et al. (2019); Zhou et al. (2020); Bejan et al. (2024); Koh (2017); Bouland et al. (2018); Zhang and Gu (2024). It is worth emphasizing that a key difference between these two resources is that entanglement remains invariant under local unitary transformations, whereas nonstabilizerness is basis-dependent, implying that it can be altered through local operations.
@@
Thus, a natural question arises: Is there a suitable measure of the genuinely non-classical resources that integrates both entanglement and nonstabilizerness? Moreover, is this measure computationally tractable? Theoretically, such a measure would deepen our understanding of the connection between these two seemingly distinct properties Iannotti et al. (2025); Szombathy et al. (2025), while practically it would help characterize algorithmic complexity and guide the selection of initial states in quantum computations and simulations. With the existence of such a measure, it becomes pertinent to identify resourceful states Chitambar and Gour (2019). Furthermore, a more intriguing question is whether this resource can be manipulated through measurement.
@@
In this paper, we address these questions by investigating quantum non-local nonstabilizerness (NN), which quantifies the minimal amount of nonstabilizerness contained in a bipartite state after optimizing over all possible local unitary transformations on each subsystem Cao et al. (2024); Cao (2024); Cepollaro et al. (2024). We first demonstrate that NN is a promising candidate for characterizing quantum resources and is applicable to any faithful and additive nonstabilizerness measure. Although previous studies have focused on providing bounds and estimates of NN, we show that NN is analytically tractable for a two-qubit pure state and directly related to the entanglement spectrum. We then investigate which states exhibit nonzero NN. Our results indicate that NN is ubiquitous in typical two-qubit pure states, with Haar-random states having the highest probability of attaining maximal NN. Extending the definition to the mixed state, we further examine the behavior of two-point NN in critical many-body states, using both the transverse-field Ising model (TFIM) at the critical point and a monitored Haar-random unitary circuit (MHC) with critical measurement rate as representative examples Li et al. (2019, 2021); Zabalo et al. (2020). We find that two-point NN decays following a power law, mirroring the behavior of mutual information. These findings establishes NN as an additional, effective probe for phase transitions from a quantum information perspective. Finally, we explore whether NN can be manipulated through measurements, akin to the manipulation of entanglement. Intriguingly, measurement-induced NN in a critical MHC decays more slowly than any correlation present before measurement, revealing a “nonstabilizerness swapping” phenomenon reminiscent of entanglement swapping Audenaert and Plenio (2005); Bennett et al. (1993).
II Non-local nonstabilizerness
@@
We consider a bipartite pure state $\left|\Psi_{AB}\right\rangle$, where $A$ and $B$ represent two subsystems. As illustrated in Fig. 1(a), the NN is defined as Cao et al. (2024); Cao (2024):
@@
where $M$ is a measure of nonstabilizerness, and $U_{A}$ and $U_{B}$ are unitary gates acting on subsystems $A$ and $B$, respectively. We will omit the subscript of $\mathcal{M}_{AB}$ when the bipartition is clear from the context. In operational terms, a finite NN indicates that the state cannot be efficiently represented as a stabilizer state, regardless of the local unitaries applied to subsystems $A$ and $B$ in an effort to reduce the nonstabilizerness. By definition, $\mathcal{M}$ is symmetric under the exchange of subsystems $A$ and $B$ and is stable under local basis rotations, which can be regarded as free operations. Since $\mathcal{M}$ can be directly related to the flatness of the entanglement spectrum, a nonzero value of $\mathcal{M}$ directly implies that the state possesses both entanglement and nonstabilizerness Tirrito et al. (2024). For any faithful nonstabilizerness measure $M$, $\mathcal{M}$ would be further guaranteed to be non-negative. Moreover, if $M$ is additive, i.e., $M(\left|\psi\right\rangle\otimes\left|\phi\right\rangle)=M(\left|\psi\right\rangle)+M(\left|\phi\right\rangle)$, then the NN is sub-additive. Specifically, for two bipartite pure states $\left|\Psi_{A_{1}B_{1}}\right\rangle$ and $\left|\Phi_{A_{2}B_{2}}\right\rangle$, we have $\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle)\geq\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. The proof can be found in Appendix A. In summary, for any additive and faithful nonstabilizerness measure, the NN is stable, sub-additive, symmetric and vanishes for both product states and stabilizer states. Given these desirable properties and clear physical significance, $\mathcal{M}$ serves as an effective measure to characterize the novel resources of a quantum state that account for both nonstabilizerness and entanglement.
@@
Although various nonstabilizerness measures satisfy faithfulness and additivity, we select $M$ to be the second stabilizer Rényi entropy (SRE) in the following, as it is well defined for qubits and avoids any minimization procedure in its calculation Leone et al. (2022); Haug et al. (2024); Lami and Collura (2023); Leone et al. (2024); Leone and Bittel (2024); Niroula et al. (2024); Odavić et al. (2023); Oliviero et al. (2022a, b); Rattacaso et al. (2023); Tarabunga et al. (2024, 2023); Frau et al. (2024); López and Kos (2024); Li et al. (2024); Fux et al. (2024); Ding et al. (2025). It is given by:
$M(\rho)=-\text{ln}\left(\frac{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{4}}{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{2}}\right),$ (2)
@@
which for pure states reduces to $M(\left|\psi\right\rangle)=-\text{ln}\sum_{P\in\mathbb{P}_{N}}(1/2^{N})|\left\langle\psi|P|\psi\right\rangle|^{4}$, where $N$ is the number of qubits and $\mathbb{P}_{N}$ represents the set of $N$-fold tensor products of Pauli matrices, denoted as $\sigma_{\mu}$ with $\mu=0,x,y,z$. It is also instructive to compare $\mathcal{M}$ with the mutual SRE introduced in previous work, which is defined as $I_{AB}(\left|\Psi_{AB}\right\rangle)=M(\left|\Psi_{AB}\right\rangle)-M(\rho_{A})-M(\rho_{B})$, where $\rho_{A}$ and $\rho_{B}$ are reduced states of subsystems $A$ and $B$, respectively Tarabunga et al. (2023); López and Kos (2024); Frau et al. (2024). While the mutual SRE captures certain aspects of the non-locality of a state’s nonstabilizerness, it lacks stability under local unitary transformations Fliss (2021); Haug and Piroli (2023), can take negative values in some cases, and generally does not align with the NN $\mathcal{M}$, as we will demonstrate later.
III Two-qubit NN
@@
In general, computing $\mathcal{M}$ is highly complicated because it involves minimizing over the entire spectrum of local unitary transformations. However, for the two-qubit case, we demonstrate that an analytic expression for $\mathcal{M}$ can be derived. We first prove the following theorem.
@@
Theorem: For a two-qubit pure state of the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$, $\mathcal{M}(\left|\psi\right\rangle)=M(\left|\psi\right\rangle)=\text{ln}(\frac{8}{7+\text{cos}(8\theta)})$.
@@
The proof is provided in Appendix B, where the key idea is that the Pauli basis representation of this specific state is in a canonical form, with the reduced density matrices of both qubits exhibiting fully polarized Bloch vectors, and the full state having a diagonal correlation matrix F. Verstraete (2002); Jevtic et al. (2014). The following important corollary arises:
@@
Corollary: For any two-qubit pure state $\left|\Psi\right\rangle$ with the bipartite entanglement spectrum $\{\text{cos}^{2}(\theta),\text{sin}^{2}(\theta)\}$, $\mathcal{M}(\left|\Psi\right\rangle)=\text{ln}(\frac{8}{7+\text{cos}(8\theta)})$.
@@
This holds because any two-qubit pure state can be transformed into the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$ by applying an appropriate local basis rotation, which can be determined via singular value decomposition. As a result, the NN is directly related to the entanglement spectrum and can be efficiently computed. It should be noted that NN has previously been understood to be approximated by quantities derived from the entanglement spectrum Cao et al. (2024). Our results show that this approximation is, in fact, exact for the two-qubit case. A direct implication is that any two-qubit pure state can have at most $\text{log}(4/3)$ amount of NN, which equals the nonstabilizerness of the H-type magic state Bravyi and Kitaev (2005). Additionally, it is noteworthy that, when combined with the sub-additivity of $\mathcal{M}$, if a many-body state can be decomposed into product states of multiple two-qubit states of the above form, an upper bound for its NN can be easily determined.
@@
Subsequently, we investigate the typical amount of NN in a random two-qubit pure state. We generate random pure states via a Haar-random two-qubit unitary gate and examine the distribution of $\mathcal{M}$ across these states. The results, shown in Fig. 1(c), indicate that the state most commonly hosts the maximal $\mathcal{M}$. While the analytical expression for this probability distribution can be derived in Appendix C, a closed-form expression for the average value of $\mathcal{M}$, analogous to that for entanglement Page (1993); Sen (1996), remains unclear. Nevertheless, numerical results suggest that the average $\mathcal{M}$ is approximately $0.192$.
IV Mixed state generalization
@@
To explore NN in a many-body state, it is first essential to extend the definition of NN to mixed states, with particular attention to the two-qubit case. For mixed states, the NN can be defined analogously to Eq. (1) as $\mathcal{M}_{AB}(\rho_{AB})=\min_{U_{A}\otimes U_{B}}M((U_{A}\otimes U_{B})\rho_{AB}(U_{A}\otimes U_{B})^{\dagger})$. Under this definition, the NN retains the properties of stability, symmetry and subadditivity for any additive measure of nonstabilizerness. By continuing to employ the second SRE, which is faithful when the mixed stabilizer states are regarded resource-free mix , the NN also vanishes for the mixed stabilizer states Leone et al. (2022); Audenaert and Plenio (2005). However, a crucial difference is that the NN does not vanish for certain separable states. This difference arises because, even for a single-qubit mixed state, such as $\rho=\frac{1}{2}(\sigma_{0}+\frac{1}{2}\sigma_{x})$, no unitary transformation can bring it back to a mixed stabilizer state. We also explore other possible nonstabilizerness measures for the mixed state, while the conclusion remains the same (see Appendix F).
@@
Focusing on the two-qubit case, it follows directly from the proof of the previous theorem that if a two-qubit mixed state $\rho$ has a Pauli basis representation in canonical form, we would have $\mathcal{M}(\rho)=M(\rho)$. However, it is important to note that not all mixed states can be transformed into the canonical form using only local unitary transformations F. Verstraete (2002). In such cases, numerical optimization is inevitable, as examples will be shown in the next section. Consider the Werner state as an analytically tractable example, which is defined as $\rho_{W}=x\left|\psi_{-}\right\rangle\left\langle\psi_{-}\right|+(1-x)\mathbb{I}/4$, with $|\psi_{-}\rangle\equiv(|01\rangle-|10\rangle)/\sqrt{2}$ Werner (1989). This state has vanishing Bloch vectors for each qubit and a diagonal correlation matrix $T$ with diagonal entries given by $-x/4$. We calculate NN and logarithmic negativity $E_{N}$, a measure of separability, with the results shown in Fig. 1(d). Although the Werner state remains separable for $x<1/3$, the NN takes on a finite value for small $x$. As $x$ approaches $1$, the state becomes more entangled, yet the NN approaches zero as the state becomes a stabilizer state. The difference between $\mathcal{M}$ and total nonstabilizerness $M$ again becomes evident if we further apply a T-gate on the first qubit. While the NN remain unaffected, the total nonstabilizerness $M$ would change significantly.